TheMexicanUSA:

Algebra. Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received: x − 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

Shadow:

Put 125,000 in for y then solve for x

Shadow:

Do you know how to do that?

TheMexicanUSA:

no, I dont know how to do that

Shadow:

\[x - 62,500 = 0.25y \rightarrow x - 62,500 = 0.25(125,000)\]

Shadow:

Multiply the two numbers on the right side (just doing orders of operations). Then use the addition property of equality to add 62,500 to both sides. This maintains the integrity of the equation, keeping it as a true statement. After you add 62,500 to the product of 0.25 times 125,000, you should see what you got.

Shadow:

Let me know if that makes sense.

TheMexicanUSA:

um is hard for me

TheMexicanUSA:

I just need the answer thats it

Shadow:

\[x - 62,500 = 0.25y\] We add in 125,000 for y, as the total # of votes cast We get: \[x - 62,500 = 0.25(125,000)\] Multiply 0.25 by 125,000 \[x - 62,500 = 31,250\] Add 62,500 to both sides, then you have your answer.

TheMexicanUSA:

hm ok Thank you very much

TheMexicanUSA:

oh yeah one thing, How do I put another question here???

Shadow:

Close this one first then you can post a new one.

TheMexicanUSA:

oh ok thanks